OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (anonymous):

Since its parallel to the line 2x+3y=5 then its slope is equal to the other line 3y=-2x+5 y=-2/3x+5 (slope = -2/3) Use the Law : \[\huge \frac{y_{2}-y_{1}}{x_{2}-x_{1}}=m\] where m = slope=-2/3 (9,-3)=(x1,y1) _______________ Need more Help ,Check this out then : http://openstudy.com/study?login#/updates/4ff5ff4de4b01c7be8c90c78